1. Find all matrices of form X = Y_120 such that x2 + 3x + 21 = 0.
Answer #1

Solution - x2+3x +21 =0 [*]**[68] D90)69]*[8] )*65 %) :) pm 43042 0 Lyntay+3y 2²+32+2 = 0 n2+32+2=0 yn+2y+3y = 0 fi 22+32 +2(2+2) (2H) =0 n=-2 and n= -1 Take (in Yurzy+3y=0 y (n+2+3)=0 y=0 or nt2+3=0 I (1) So yzo it is clear for any value of (n+2+3)